OpenStudy (anonymous):

Solve the inequality and then graph its solution:

OpenStudy (anonymous):

\[5\le -\frac{ s }{ 2 } +4\]

OpenStudy (anonymous):

got a clue

OpenStudy (anonymous):

just one at the most

OpenStudy (anonymous):

i am running out of time

OpenStudy (anonymous):

>:-(

OpenStudy (anonymous):

come on

OpenStudy (anonymous):

say some thing

OpenStudy (anonymous):

>:-(

OpenStudy (anonymous):

17:33 is left on the clock

OpenStudy (anonymous):

\[a. s \ge-2\]

OpenStudy (anonymous):

\[b.s \le-2\]

OpenStudy (anonymous):

ais 2

OpenStudy (anonymous):

\[c.s \ge-2\]

OpenStudy (anonymous):

\[d.s \le2\]

OpenStudy (anonymous):

now can you find it out

OpenStudy (anonymous):

if your steal there